# How Do Developers Interact with AI? An Exploratory Study on Modeling Developer Programming Behavior

### 28 Mar 2026 v1

@@ 5. RQ2: Aggregated Behavioral Patterns
To address RQ2, our analysis focused on identifying the differences between AI-assisted and non-AI participants. To make the interpretation easier, we visualize our results with explanations. The bar charts in Figure 3, Figure 4, and Figure 5 present the respective results for the aggregated behavioral patterns that emerged from intention, action, supporting tool, and emotion dimensions.
For each category, we first reported its occurrence frequency by collecting the number of occurrences and calculating the mean number of occurrences for each participant. We computed the relative difference per task and averaged them. A positive value means the category is more (↑) frequent in the AI groups, and a negative value means less (↓). Besides, we discuss patterns with statistically significant differences ($\alpha=0.05$). Specifically, for the categorical dimensions, we computed per-participant rates by dividing the number of intervals annotated with a given category by the total intervals per participant. For emotion, we calculated each participant’s mean valence score. We then fitted ordinary least squares (OLS) regression (Cohen et al., 2013) with heteroscedasticity-robust standard errors (HC3) (Long and Ervin, 2000). All models controlled for AI condition and task type. For the categorical dimensions, we applied Benjamini–Hochberg false discovery rate (FDR) correction (Benjamini and Hochberg, 1995) to account for multiple comparisons, and primarily discuss categories that remained statistically significant after correction.
@@
$\bullet$ Intention. We found that 6 of the 12 intentions occurred more frequently among AI participants than among non-AI participants. They are: To Verify Existing Code, To Resolve Errors, To Optimize Code, To Implement Code, To Evaluate Suggestions, and To Design/Plan New Code.
@@
Specifically, AI-assisted participants formed the intention To Evaluate Suggestions significantly more frequently ($\beta=0.069,p<0.001$). This suggests that using AI assistants often leads to occurrences of participants seeking to assess and evaluate their code. Other intentions, such as To Resolve Errors, To Implement Code, and To Design/Plan New Code, showed directional differences between AI-assisted and non-AI groups but did not reach statistical significance after correction. These patterns therefore suggest descriptive trends that programming with AI keeps participants focusing on actively “creating” code (i.e., with AI assistance), such as designing, evaluating, and resolving these generated results (e.g., To Resolve Errors ($\beta=0.0748$)) .
@@
$\bullet$ Action. Among the 16 actions, we found that 9 of them occurred more frequently with AI-assisted participants. After statistical testing and multiple-comparison correction, a subset of actions exhibited statistically significant differences. Particularly, Executing the Code ($\beta=0.096,p<0.001$) and Editing Query/Prompt ($\beta=0.028,p<0.001$) occurred significantly more frequently. This can be explained by the identified vibe coding phenomenon (Wikipedia, 2025; Sarkar and Drosos, 2025). Interestingly, with this phenomenon, we found that they performed Reading and Comprehension significantly less frequently ($\beta=-0.146,p<0.001$). These results suggest that AI assistants transform participants’ programming experience into an iterative cycle of prompting and rapid verification (trial-and-error), reducing the time spent on deep manual code reading and environment setup.
For example, when AI-assisted participants implemented PyT on querying GitHub’s API service, they often directly copy-pasted the AI-generated code and then performed quick verification by executing the code without taking specific actions to try and understand the code. If the code functioned as expected, they directly progressed to the next step of the task.
This pattern suggests that participants often perceived AI outputs as sufficiently usable to proceed without deep inspection. The result is also consistent with our findings in terms of task completion rates, i.e., the AI-assisted group produced a higher completeness ratio than the non-AI group (described in Section 3).
@@
$\bullet$ Supporting Tool. We observed that AI-assisted developers relied significantly less on Search Engines ($\beta=-0.246,p<0.001$). This is in line with the decline in internet traffic to Stack Overflow since the release of ChatGPT (Overflow, 2023), and suggests that AI assistants become the primary knowledge interface. Considered together with the action-level results, this pattern indicates that AI assistance promotes an iterative “prompt–run–check” practice, reducing time spent on deep manual reading and external information seeking. Under this practice, tool use is primarily oriented toward rapid execution and verification rather than extended web searching.
@@
$\bullet$ Emotion. We found that feelings of Positive, Slightly Positive, and Neutral were more commonly expressed by AI-assisted participants, while feelings of Negative and Slightly Negative were more common for non-AI-assisted participants. Primary OLS results showed that AI assistance was associated with significantly more positive reported emotions. The AI group reported significantly more positive emotions ($\beta=0.632,p=0.016$) compared to the non-AI group. This indicates that participants allowed to use AI assistants for the programming task experienced more positive feelings than those without AI.
@@
To assess the robustness of our findings, we conducted an interval-level analysis using mixed-effects models (LMM) (Laird and Ware, 1982) for emotion and generalized estimating equations (GEE) (Liang and Zeger, 1986) for categorical dimensions, accounting for participant-level clustering. Across the Intention dimension, the same core categories remained significant with consistent effect directions. For Action, all categories significant in the robustness check were also identified in the primary analysis, while the primary analysis identified a small number of additional action categories. For Supporting Tools, the robustness check identified Proprietary Developer Platform as an additional significant category. Importantly, no statistically significant findings in the primary analysis were contradicted by the robustness checks. All detailed results are provided in our replication package (Wu et al., 2025).
@@ 6. RQ3: Sequential Behavioral Patterns
To answer our RQ3, we conducted a sequential pattern mining analysis with a three-step, mixed-methods procedure: (1) Automated pattern mining: For each S-IASE dimension, we applied CloSpan, a closed sequential pattern mining algorithm, based on the annotated AI-assisted intervals produced by 62 of the 76 participants. We chose CloSpan because it enumerates closed patterns via a prefix-projected depth-first search strategy, producing a non-redundant set of frequent sequences (Yan et al., 2003). Prior studies have applied closed (or frequent) sequential pattern mining to developer data (e.g., IDE usage mining (Damevski et al., 2016), and API usage (Xie and Pei, 2006)), demonstrating its suitability for analyzing developer behavioral data. We applied the CloSpan algorithm with a minimum support threshold of 10%, in-line with prior work (Damevski et al., 2016), to the four dimensions of our S-IASE model: Intention, Action and Support Tool Pair(s), and Emotion. This step provided a non-redundant set of frequent candidate sequential patterns. (2) Manual interpretation and selection of patterns: Two researchers independently inspected the frequent patterns produced by Step 1. For each pattern, they: (i) examined its occurrences in the underlying participant sequences and screen-recording context, (ii) assessed whether the pattern represented a coherent, interpretable behavioral regularity (e.g., the “Trust but Verify” sequence To Understand Context → To Implement Code → To Design/Plan New Code), and (iii) grouped semantically similar patterns. Disagreements about inclusion or grouping were resolved through discussion until consensus was reached. This step reduced the raw mined output to a smaller set of behaviorally meaningful patterns that we considered for reporting. (3) Triangulation with interview data and existing knowledge: We then triangulated these candidate patterns with our 14 follow-up interviews (analyzed using STGT) and existing concepts in prior literature (e.g., vibe coding (Sarkar and Drosos, 2025; Wikipedia, 2025), trust in AI assistants (Choudhuri et al., 2024), and imposter phenomenon (Guenes et al., 2024)). Patterns were retained only when supported by both sequential data and qualitative evidence. Through this process, the sequential patterns are grounded in both quantitative evidence from the sequence mining algorithm and qualitative evidence from interviews and prior knowledge.
@@ 6.1. Trust but Verify: Navigating Confidence and Skepticism in AI-Assisted Coding
This pattern is particularly noteworthy because it reverses the traditional “Design-then-Implement” paradigm. With AI assistance, developers often leverage AI to implement code first to gain a concrete foothold in the codebase, and then proceed to refine the architecture or design (i.e., To Design/Plan New Code) based on the AI’s output. We also found that the willingness to adopt an AI assistant is influenced by task familiarity.
Prior work (Barke et al., 2023) indicates people are more willing to adopt AI assistants for familiar tasks; we extend this by finding that developers were likely to accept AI-generated suggestions when they perceived a task as either highly familiar or highly unfamiliar. For tasks that participants were familiar with but were not fluent in developing (e.g., setting up and establishing a MySQL Database connection), participants exhibited skepticism towards AI assistants despite correct suggestions. They often hesitated to trust AI-generated suggestions without external validation. However, P6 did not immediately trust the link ChatGPT generated and instead Googled for the result, despite ChatGPT providing the correct link.
@@
I usually take what Copilot gives me and try it out. If it works, I keep it; if not, I tweak it or ask ChatGPT to refine it.(P7) This back-and-forth process showed that developers are not passively accepting AI-generated code; they remain actively engaged, evaluating and modifying suggestions as needed.
@@ 6.2.1. The Tension Between Speed and Understanding
While AI assistance resulted in higher emotional stability and faster completion times, many AI-assisted participants reported guilty feelings. The set time limit of 25 minutes nudged participants to prioritize speed over understanding. A consequence of prioritizing speed and completion of the tasks was that it left some participants feeling conflicted about the depth of their learning:
I felt bad that I just completed it, but I didn’t learn more. If I want to use the REST API in the future, I’ll need to ask ChatGPT again.(P6)
@@ 6.2.2. Impostor Phenomenon and Self-Criticism
This emotional tension was further compounded by manifestations of the impostor phenomenon, which is the belief that one’s success is undeserved or luck rather than competence. Researchers have found impostor phenomenon to negatively impact perceived productivity in software engineers (Guenes et al., 2024).
@@
This self-criticism carried over into debugging and troubleshooting. When P2 encountered a problem with AI-generated code, P2’s first instinct was to assume that the fault lay with their machine rather than the AI.
My first thought was, “Is my laptop broken?” I wasted 3 minutes checking my laptop.(P2)
@@ 6.3. Show Me, Don’t Tell Me: The Need for Visual Clarity in AI Assistants
The intention to obtain required knowledge was important because this sequence emerged as a notable finding during follow-up interviews. Participants reported that this intention was time-consuming, especially when dealing with procedural tasks such as configuring API tokens or looking up the related documentation. Although standard AI assistant prompting is straightforward, participants found text-only AI suggestions insufficient in providing in-depth explanations. For example, when ChatGPT provided written instructions for GitHub API setup, developers like P1 and P4 abandoned the assistant altogether. They turned to platforms like Stack Overflow or GitHub’s official documentation to provide step-by-step instructions.
ChatGPT told me to “add the token” but where exactly? some screenshots or visual steps would’ve saved me 10 minutes of guessing.(P1)
@@
This pattern highlights a modality mismatch: when a developer’s Intention is To Set Up Project, the purely text-based modality of current AI assistants fails, forcing a fallback to traditional Search Engines to find diagrams or visual documentation. This pattern reflects developers’ need for concrete visual guidance versus abstract textual explanations. P4 expressed greater trust in answers when supported by visual aids:
@@
Similarly, P7 spent a significant amount of the study time on the database setup and connection for the Java assignment. Despite frequently prompting an AI assistant to debug the database setup and connection errors, these kinds of tasks were more complex than simply copying and pasting a few lines of code. The AI assistant provided various instructional sets, but could not visually demonstrate to the participant the correct order of steps to take to achieve the goal. This forced participants to repeatedly interpret textual suggestions, resulting in high cognitive load and time costs. Future work from researchers should investigate how to help AI assistant practitioners add features that support visual guidance.
@@ 7.1. Situating S-IASE in the Landscape of Developer–AI Interaction Research
There are significant differences between our proposed S-IASE model and the prior work on developer-AI interaction. For example, while Google’s SIA model (Research, 2023) describes coding states primarily through intentions and actions, our model incorporates explicit dimensions of supporting tool and emotion. This addition essentially highlights how affective states (i.e., emotions) intertwine with cognitive processes (i.e., actions) during programming, complementing and refining the abstractions used in earlier models such as SIA (Research, 2023) and CUPS (Mozannar et al., 2024).
In doing so, our work challenges the view of developer states as a single coding activity, showing instead that intentions, actions, tools, and emotions evolve together as developers interact and collaborate with AI assistants.
@@ 7.2. Implications for Researchers
$\bullet$ Model for developer-AI programming behavior: The rapid growth and adoption of AI assistants in software development have placed greater emphasis on understanding how developers use AI assistants. The model proposed in our work is a step towards a deeper understanding of all the intricacies involved in developer behavior when they interact with AI assistants for coding. Future work could explore some of the behavior patterns uncovered in more depth or incorporate additional dimensions to enhance the model.
$\bullet$ Trusting AI assistants: Our findings indicate developers are more willing to use AI assistants when they are highly familiar or completely unfamiliar with a task. This differs from previous literature on factors impacting developer trust of AI suggestions, such as quality of the suggestion and context (Brown et al., 2024). Further research should focus on deriving deeper insights about what influences developer trust in AI-generated answers, and more importantly, exploring how AI assistants should be enhanced to improve developer trust.
$\bullet$ Guilt when using AI: Impostor syndrome has been documented to occur in software engineering and negatively impact perceived productivity (Guenes et al., 2024). We found developers self-criticizing during the use of AI assistants for programming and blaming themselves for over-relying on AI. Future research would benefit from empirical research on the extent of the impostor phenomenon from developers use of AI assistants. Future research should aim to uncover strategies to mitigate the negative emotions from using AI.

### 26 Apr 2026 v2

@@ 5. RQ2: Aggregated Behavioral Patterns
To address RQ2, our analysis focused on identifying the differences between AI-assisted and non-AI participants. To make the interpretation easier, we visualize our results with explanations. The bar charts in Figure 3, Figure 4, and Figure 5 present the respective results for the aggregated behavioral patterns that emerged across intention, action, supporting tool, and emotion dimensions. For each category, we first report its occurrence frequency by collecting the number of occurrences and calculating the mean number of occurrences for each participant. We computed the relative difference per task and averaged them. A positive value means the category is more (↑) frequent in the AI groups, and a negative value means less (↓). Besides, we discuss patterns with statistically significant differences ($\alpha=0.05$). Specifically, for the categorical dimensions, we computed per-participant rates by dividing the number of intervals annotated with a given category by the total intervals per participant. For emotion, we calculated each participant’s mean valence score. We then fitted ordinary least squares (OLS) regression (Cohen et al., 2013) with heteroscedasticity-robust standard errors (HC3) (Long and Ervin, 2000). All models controlled for AI condition and task type. For the categorical dimensions, we applied Benjamini–Hochberg false discovery rate (FDR) correction (Benjamini and Hochberg, 1995) to account for multiple comparisons, and primarily discuss categories that remained statistically significant after correction.
@@
$\bullet$ Intention. We found that 6 of the 12 intentions occurred more frequently among AI participants than among non-AI participants. They are: To Verify Existing Code, To Resolve Errors, To Optimize Code, To Implement Code, To Evaluate Suggestions, and To Design/Plan New Code. Specifically, AI-assisted participants formed the intention To Evaluate Suggestions significantly more frequently ($\beta=0.069,p<0.001$). This suggests that using AI assistants often leads to occurrences of participants seeking to assess and evaluate their code. Other intentions, such as To Resolve Errors, To Implement Code, and To Design/Plan New Code, showed directional differences between AI-assisted and non-AI groups but did not reach statistical significance after correction. These patterns therefore suggest descriptive trends that programming with AI keeps participants focusing on actively “creating” code (i.e., with AI assistance), such as designing, evaluating, and resolving the generated results (e.g., To Resolve Errors ($\beta=0.0748$)).
@@
$\bullet$ Action. Among the 16 actions, we found that 9 of them occurred more frequently with AI-assisted participants. After statistical testing and multiple-comparison correction, a subset of actions exhibited statistically significant differences. Particularly, Executing the Code ($\beta=0.096,p<0.001$) and Editing Query/Prompt ($\beta=0.028,p<0.001$) occurred significantly more frequently. This can be explained by the identified vibe coding phenomenon (Wikipedia, 2025; Sarkar and Drosos, 2025). Interestingly, with this phenomenon, we found that they performed Reading and Comprehension significantly less frequently ($\beta=-0.146,p<0.001$). These results suggest that AI assistants transform participants’ programming experience into an iterative cycle of prompting and rapid verification (trial-and-error), reducing the time spent on deep manual code reading and environment setup. For example, when AI-assisted participants implemented PyT on querying GitHub’s API service, they often directly copy-pasted the AI-generated code and then performed quick verification by executing the code without taking specific actions to try and understand the code. If the code functioned as expected, they directly progressed to the next step of the task. This pattern suggests that participants often perceived AI outputs as sufficiently usable to proceed without deep inspection. This result is also consistent with our findings in terms of task completion rates, i.e., the AI-assisted group achieved a higher completion rate than the non-AI group (described in Section 3).
@@
$\bullet$ Supporting Tool. We observed that AI-assisted developers relied significantly less on Search Engines ($\beta=-0.246,p<0.001$). This is in line with the decline in internet traffic to Stack Overflow since the release of ChatGPT (Overflow, 2023), and suggests that AI assistants have become the primary knowledge interface. Together with the action-level results, this pattern indicates that AI assistance promotes an iterative “prompt–run–check” practice, reducing time spent on deep manual reading and external information seeking. Under this practice, tool use is primarily oriented toward rapid execution and verification rather than extended web searching.
@@
$\bullet$ Emotion. We found that Positive, Slightly Positive, and Neutral emotions were more commonly expressed by AI-assisted participants, while Negative and Slightly Negative emotions were more common for non-AI-assisted participants. Primary OLS results showed that AI assistance was associated with significantly more positive reported emotions. The AI group reported significantly more positive emotions ($\beta=0.632,p=0.016$) compared to the non-AI group. This indicates that participants allowed to use AI assistants for the programming task experienced more positive feelings than those without AI.
@@
To assess the robustness of our findings, we conducted an interval-level analysis using mixed-effects models (LMM) (Laird and Ware, 1982) for emotion and generalized estimating equations (GEE) (Liang and Zeger, 1986) for categorical dimensions, accounting for participant-level clustering. Across the Intention dimension, the same core categories remained significant with consistent effect directions. For Action, all categories significant in the robustness check were also identified in the primary analysis, while the primary analysis identified a small number of additional action categories. For Supporting Tools, the robustness check identified Proprietary Developer Platform as an additional significant category. Importantly, no statistically significant findings in the primary analysis were contradicted by the robustness checks. All detailed results are provided in our replication package (Wu et al., 2026).
@@ 6. RQ3: Sequential Behavioral Patterns
To answer RQ3, we conducted a sequential pattern mining analysis with a three-step, mixed-methods procedure: (1) Automated pattern mining: For each S-IASE dimension, we applied CloSpan, a closed sequential pattern mining algorithm, based on the annotated AI-assisted intervals produced by 62 of the 76 participants. We chose CloSpan because it enumerates closed patterns via a prefix-projected depth-first search strategy, producing a non-redundant set of frequent sequences (Yan et al., 2003). Prior studies have applied closed (or frequent) sequential pattern mining to developer data (e.g., IDE usage mining (Damevski et al., 2016) and API usage (Xie and Pei, 2006)), demonstrating its suitability for analyzing developer behavioral data. We applied the CloSpan algorithm with a minimum support threshold of 10%, in line with prior work (Damevski et al., 2016), to the four dimensions of our S-IASE model: Intention, Action and Supporting Tool Pair(s), and Emotion. This step provided a non-redundant set of frequent candidate sequential patterns. (2) Manual interpretation and selection of patterns: Two researchers independently inspected the frequent patterns produced by Step 1. For each pattern, they: (i) examined its occurrences in the underlying participant sequences and screen-recording context, (ii) assessed whether the pattern represented a coherent, interpretable behavioral regularity (e.g., the “Trust but Verify” sequence To Understand Context → To Implement Code → To Design/Plan New Code), and (iii) grouped semantically similar patterns. Disagreements about inclusion or grouping were resolved through discussion until consensus was reached. This step reduced the raw mined output to a smaller set of behaviorally meaningful patterns that we considered for reporting. (3) Triangulation with interview data and existing knowledge: We then triangulated these candidate patterns with our 14 follow-up interviews (analyzed using STGT) and existing concepts in prior literature (e.g., vibe coding (Sarkar and Drosos, 2025; Wikipedia, 2025), trust in AI assistants (Choudhuri et al., 2025), and impostor phenomenon (Guenes et al., 2024)). Patterns were retained only when supported by both sequential data and qualitative evidence. Through this process, the sequential patterns are grounded in both quantitative evidence from the sequence mining algorithm and qualitative evidence from interviews and prior knowledge.
@@ 6.1. Trust but Verify: Navigating Confidence and Skepticism in AI-Assisted Coding
This pattern is particularly noteworthy because it reverses the traditional “Design-then-Implement” paradigm. With AI assistance, developers often leverage AI to implement code first to gain a concrete foothold in the codebase, and then proceed to refine the architecture or design (i.e., To Design/Plan New Code) based on the AI’s output. We also found that the willingness to adopt an AI assistant is influenced by task familiarity. Prior work (Barke et al., 2023) indicates that people are more willing to adopt AI assistants for familiar tasks; we extend this by finding that developers were more likely to accept AI-generated suggestions when they perceived a task as either highly familiar or highly unfamiliar. For tasks that participants were familiar with but were not fluent in developing (e.g., setting up and establishing a MySQL Database connection), participants exhibited skepticism toward AI assistants despite correct suggestions. They often hesitated to trust AI-generated suggestions without external validation. For example, P6 did not immediately trust the link ChatGPT generated and instead Googled for the result, despite ChatGPT providing the correct link.
@@
I usually take what Copilot gives me and try it out. If it works, I keep it; if not, I tweak it or ask ChatGPT to refine it.(P7)
This back-and-forth process showed that developers are not passively accepting AI-generated code; they remain actively engaged, evaluating and modifying suggestions as needed.
@@ 6.2.1. The Tension Between Speed and Understanding
While AI assistance resulted in higher emotional stability and faster completion times, many AI-assisted participants reported guilty feelings. The 25-minute time limit nudged participants to prioritize speed over understanding. A consequence of prioritizing speed and completion of the tasks was that it left some participants feeling conflicted about the depth of their learning:
@@ 6.2.2. Impostor Phenomenon and Self-Criticism
This emotional tension was further compounded by manifestations of the impostor phenomenon, which is the belief that one’s success is undeserved or due to luck rather than competence. Researchers have found impostor phenomenon to negatively impact perceived productivity in software engineers (Guenes et al., 2024).
@@
This self-criticism carried over into debugging and troubleshooting. When P2 encountered a problem with AI-generated code, their first instinct was to assume that the fault lay with their machine rather than the AI.
My first thought was, ‘Is my laptop broken?’ I wasted 3 minutes checking my laptop.(P2)
@@ 6.3. Show Me, Don’t Tell Me: The Need for Visual Clarity in AI Assistants
The intention To Obtain Required Knowledge was important because this sequence emerged as a notable finding during follow-up interviews. Participants reported that this intention was time-consuming, especially when dealing with procedural tasks such as configuring API tokens or looking up the related documentation. Although standard AI assistant prompting is straightforward, participants found text-only AI suggestions insufficient for providing in-depth explanations. For example, when ChatGPT provided written instructions for GitHub API setup, developers like P1 and P4 abandoned the assistant altogether. They turned to platforms like Stack Overflow or GitHub’s official documentation to find step-by-step instructions.
ChatGPT told me to ‘add the token’ but where exactly? Some screenshots or visual steps would’ve saved me 10 minutes of guessing.(P1)
@@
This pattern highlights a modality mismatch: when a developer’s Intention is To Set Up Project Environment, the purely text-based modality of current AI assistants fails, forcing a fallback to traditional Search Engines to find diagrams or visual documentation. This pattern reflects developers’ need for concrete visual guidance versus abstract textual explanations. P4 expressed greater trust in answers when supported by visual aids:
@@
Similarly, P7 spent a lot of time on the database setup and connection for the Java assignment. Despite frequently prompting an AI assistant to debug the database setup and connection errors, these kinds of tasks were more complex than simply copying and pasting a few lines of code. The AI assistant provided various instructional sets, but could not visually demonstrate to the participant the correct order of steps to take to achieve the goal. This forced participants to repeatedly interpret textual suggestions, resulting in high cognitive load and time costs. Future work should investigate how to help developers of AI assistants add features that support visual guidance.
@@ 7.1. Situating S-IASE in the Landscape of Developer–AI Interaction Research
There are significant differences between our proposed S-IASE model and prior work on developer-AI interaction. For example, while Google’s SIA model (Research, 2023) describes coding states primarily through intention and action, our model incorporates explicit dimensions of supporting tool and emotion. This addition essentially highlights how affective states (i.e., emotions) intertwine with cognitive and behavioral processes (i.e., actions) during programming, complementing and refining the abstractions used in earlier models such as SIA (Research, 2023) and CUPS (Mozannar et al., 2024). In doing so, our work challenges the view of developer states as a single coding activity, showing instead that intentions, actions, tools, and emotions evolve together as developers interact and collaborate with AI assistants.
@@ 7.2. Implications for Researchers
$\bullet$ Model for developer-AI programming behavior: The rapid growth and adoption of AI assistants in software development have placed greater emphasis on understanding how developers use them. The model proposed in our work is a step toward a deeper understanding of all the intricacies involved in developer behavior when they interact with AI assistants for coding. Future work could explore some of the behavior patterns uncovered in more depth or incorporate additional dimensions to enhance the model.
$\bullet$ Trusting AI assistants: Our findings indicate that developers are more willing to use AI assistants when they are highly familiar or completely unfamiliar with a task. This differs from previous literature on factors impacting developer trust in AI suggestions, such as quality of the suggestion and context (Brown et al., 2024). Further research should focus on deriving deeper insights about what influences developer trust in AI-generated answers, and more importantly, exploring how AI assistants should be enhanced to improve developer trust.
$\bullet$ Guilt when using AI: Impostor syndrome has been documented to occur in software engineering and negatively impact perceived productivity (Guenes et al., 2024). We found developers self-criticizing during the use of AI assistants for programming and blaming themselves for over-relying on AI. Future research should empirically investigate the extent of the impostor phenomenon in AI-assisted development and uncover strategies to mitigate these negative emotions from using AI.